OpenStudy (anonymous):

The average salary for computer programmmers is $7,690 less than twice the average salary for carpenters. Combined, their avearge salaries are $99,830. Determine the average salary for each job.

OpenStudy (gw2011):

Let c=carpenters salary Let cp=computer programmers salary cp=2c-7690 2c-7690+c=99,830 2c+c=99,830+7690 3c=107,520 c=$35,840 This is the carpenters salary cp=2c-7690 cp=(2)(35,840)-7690 cp=71680-7690 cp=$63,990 This is the computer programmers salary Therefore, the computer programmers salary is $63,990 and the carpenters salary is $35,840